'' I can tell that 918 273 645 can be divided by 9 without calculation '' what evidence does the speaker have to be able to make this statements?
1
Expert's answer
2020-06-29T15:04:12-0400

A number is divisible by 9 if the sum of its digits is divisible by 9 (a criterion of divisibility by 9)


So 9+1+8+2+7+3+6+4+5=9+(1+8)+(2+7)+(3+6)+(4+5)=9+9+9+9+9


Of course, this number is divisible by 9.


That's why according to the rule we can say that the number 918273645 is divisible by 9 without calculations.